# Remote Keylogging Attacks in Multi-user VR Applications

### v1

@@ Abstract
As Virtual Reality (VR) applications grow in popularity, they have bridged distances and brought users closer together. However, with this growth, there have been increasing concerns about security and privacy, especially related to the motion data used to create immersive experiences. In this study, we highlight a significant security threat in multi-user VR applications, which are applications that allow multiple users to interact with each other in the same virtual space. Specifically, we propose a remote attack that utilizes the avatar rendering information collected from an adversary’s game clients to extract user-typed secrets like credit card information, passwords, or private conversations. We do this by (1) extracting motion data from network packets, and (2) mapping motion data to keystroke entries. We conducted a user study to verify the attack’s effectiveness, in which our attack successfully inferred 97.62% of the keystrokes. Besides, we performed an additional experiment to underline that our attack is practical, confirming its effectiveness even when (1) there are multiple users in a room, and (2) the attacker cannot see the victims. Moreover, we replicated our proposed attack on four applications to demonstrate the generalizability of the attack. These results underscore the severity of the vulnerability and its potential impact on millions of VR social platform users.
@@ 1 Introduction
Our attack consists of four steps, each extracting more fine-grained information about the typing activity from the previous step. In these four steps, we aim to understand and accurately leverage the data in network packets, converting the data to recover keystrokes and thereby executing the remote keylogging attack. We evaluated our attack through a user study conducted on Rec Room [8], one of the most widely used multi-user VR applications with more than 15 million users [39]. We successfully reconstructed the typed secrets at top-1 accuracy of 97.62%. Our attack results demonstrate that we can infer almost all user-typed information correctly, even though remote settings offer lower-fidelity information compared to local settings. Furthermore, we performed an additional experiment in which (1) there are multiple users in the room, and (2) the attacker does not see any other user (from their application client’s point of view). Our attack achieves comparable performance under this setting (top-1 accuracy of 97.53%), demonstrating the practicality of our attack. Lastly, we replicated our attack on three additional applications  [^1] The names of the three additional applications evaluated will be updated following the 90-day responsible disclosure period. and performed user studies on them, in which we achieved comparable performance across all applications (top-1 accuracy of 98.24%, 98.27%, and 99.07% respectively from the additional applications).
This result further demonstrates that our attack is generalizable across applications. We reported our attack to Rec Room, the three additional applications, SteamVR, and Unity. The developers of Rec Room and SteamVR [12] acknowledged the issue and Rec Room also awarded us a bounty for the vulnerability.
@@
(5) We propose countermeasures for multi-user VR applications to mitigate our remote keylogging attacks and avoid privacy leakage associated with transmitting motion data across the Internet.
@@ 5 Data Collection
Other Multi-user VR Applications. To show that our attack generalizes beyond Rec Room, we tested it on three additional popular applications from diverse genres. Specifically, we conducted an end-to-end attack on each of these applications following the same attack procedure (Section 4) and setup of the experiment on Rec Room (the User Study Typing Trials experiment). The details for the experiment setup will be released 90 days after the disclosure.
@@ 6.5 Our Attack Generalizes Across Applications
By performing our proposed attack on the additional three applications, we have successfully recovered the motion data from the collected traffic and infer the key participants typed. For more details on the recovered packets, please refer to the examples in Appendix A.
@@ 7 Keylogging is Possible Even with Partial Reverse Engineered Packets
We have also performed further analysis of the attack results in Appendix E, focusing on inferring keystrokes with different amounts of training data, different hands, and during different typing tasks.
@@ 8 Discussion
User Awareness and Concerns. From the results of the survey (see Appendix F for the survey questions), none of the participants had ever imagined that their input in multi-user VR applications could be stolen by adversaries. However, after disclosure, 12 out of 20 participants expressed concerns regarding this type of attack (Score > 4). These findings suggest that VR users do recognize the potential harm from such attacks. The user study results highlight the importance of educating VR users about potential security and privacy threats, such as our keylogging attacks in multi-user VR applications. This education can serve as a warning, encouraging users to exercise caution and prevent privacy leakage.
@@
Impacts of Our Attack. We have reported our attack findings to all of the applications we evaluated, as well as SteamVR and Unity. Rec Room, and SteamVR responded, acknowledging the vulnerability. Furthermore, Rec Room classified it as P3 severity (Medium: Vulnerabilities that affect multiple users and require little or no user interaction to trigger). This underscores the fact that our attack is feasible and poses a privacy risk to numerous users. At the time of writing, we still have ongoing discussions with SteamVR and Unity.
@@ Appendix A Example of a Packet Parsed with Generic Photon Protocol Parser
In Figure 13, we show how the packet payload is parsed by a generic parser for the Photon protocol (Section 4.2). Photon is used to exchange objects between game clients and the Rec Room server. Note that, at this stage of parsing, we (the attacker) do not (yet) know the meaning (semantics) of specific fields of these objects, which are structured in different ways for different applications. However, the Photon parser can decode the raw packet payload into objects.
@@ Appendix D Other Results for the Keylogging Attack
Appendix E More Analysis for Attack on Partially Reverse Engineered Packets

### v2

@@ Abstract
As Virtual Reality (VR) applications grow in popularity, they have bridged distances and brought users closer together. However, with this growth, there have been increasing concerns about security and privacy, especially related to the motion data used to create immersive experiences. In this study, we highlight a significant security threat in multi-user VR applications, which are applications that allow multiple users to interact with each other in the same virtual space. Specifically, we propose a remote attack that utilizes the avatar rendering information collected from an adversary’s game clients to extract user-typed secrets like credit card information, passwords, or private conversations. We do this by (1) extracting motion data from network packets, and (2) mapping motion data to keystroke entries. We conducted a user study to verify the attack’s effectiveness, in which our attack successfully inferred 97.62% of the keystrokes. Besides, we performed an additional experiment to underline that our attack is practical, confirming its effectiveness even when (1) there are multiple users in a room, and (2) the attacker cannot see the victims. Moreover, we replicated our proposed attack on four applications to demonstrate the generalizability of the attack. Lastly, we proposed a defense against the attack, which has been implemented by major players in the VR industry. These results underscore the severity of the vulnerability and its potential impact on millions of VR social platform users.
@@ 1 Introduction
Our attack consists of four steps, each extracting more fine-grained information about the typing activity from the previous step. In these four steps, we aim to understand and accurately leverage the data in network packets, converting the data to recover keystrokes and thereby executing the remote keylogging attack. We evaluated our attack through a user study conducted on Rec Room [8], one of the most widely used multi-user VR applications with more than 15 million users [44]. We successfully reconstructed the typed secrets at top-1 accuracy of 97.62%. Our attack results demonstrate that we can infer almost all user-typed information correctly, even though remote settings offer lower-fidelity information compared to local settings. Furthermore, we performed an additional experiment in which (1) there are multiple users in the room, and (2) the attacker does not see any other user (from their application client’s point of view). Our attack achieves comparable performance under this setting (top-1 accuracy of 97.53%), demonstrating the practicality of our attack. Lastly, we replicated our attack on three additional applications and performed user studies on them, in which we achieved comparable performance across all applications (top-1 accuracy of 98.24%, 98.27%, and 99.07% respectively from the additional applications). This result further demonstrates that our attack is generalizable across applications. We reported our attack to Rec Room, the three additional applications, SteamVR, and Unity. The developers of Rec Room, Sing Together: VR Karaoke [15] and SteamVR [16] acknowledged the issue and Rec Room also awarded us a bounty for the vulnerability. Furthermore, SteamVR and Rec Room have created patches to defend against our proposed attack, following our defense suggestions as outlined in Section 8.
@@
(5) We propose countermeasures for multi-user VR applications to mitigate our remote keylogging attacks and avoid privacy leakage associated with transmitting motion data across the Internet. Rec Room and SteamVR have adopted and implemented these countermeasures.
@@ 5 Data Collection
Other Multi-user VR Applications. To show that our attack generalizes beyond Rec Room, we tested it on three additional popular applications from diverse genres: Galaxity [12], Sing Together: VR Karaoke [15], and oVRshot [13] (see Appendix E for more details). Specifically, we conducted an end-to-end attack on each of these applications following the same attack procedure (Section 4) and setup of the experiment on Rec Room (the User Study Typing Trials experiment). Among these applications, we targeted keystrokes on the full keyboards of two (Galaxity and Sing Together: VR Karaoke), whereas in oVRshot, attention was directed to the numeric keyboard, given that it is the only type available. During the typing experiment, which was conducted with three participants for each application, we collected a total of 3,423 clicks for Galaxity, 3,473 clicks for Sing Together: VR Karaoke, and 760 clicks for oVRshot (numbers trials only). Note that for passwords and sentences trials, we selected random words that can have slightly different lengths, which leads to a slight variance in total keystrokes for different applications.
@@ 6.5 Our Attack Generalizes Across Applications
By performing our proposed attack on Galaxity [12], Sing Together: VR Karaoke [15], and oVRshot [13], we have successfully recovered the motion data from the collected traffic and infer the key participants typed. For more details on the recovered packets, please refer to the examples in Appendix A.
@@ 7 Keylogging is Possible Even with Partial Reverse Engineered Packets
We have also performed further analysis of the attack results in Appendix F, focusing on inferring keystrokes with different amounts of training data, different hands, and during different typing tasks.
@@ 8 Discussion
User Awareness and Concerns. From the results of the survey (see Appendix G for the survey questions), none of the participants had ever imagined that their input in multi-user VR applications could be stolen by adversaries. However, after disclosure, 12 out of 20 participants expressed concerns regarding this type of attack (Score > 4). These findings suggest that VR users do recognize the potential harm from such attacks. The user study results highlight the importance of educating VR users about potential security and privacy threats, such as our keylogging attacks in multi-user VR applications. This education can serve as a warning, encouraging users to exercise caution and prevent privacy leakage.
@@
Impacts of Our Attack. We have reported our attack findings to all of the applications we evaluated, as well as SteamVR and Unity. Rec Room, Sing Together: VR Karaoke and SteamVR responded, acknowledging the vulnerability. Furthermore, Rec Room classified it as P3 severity (Medium: Vulnerabilities that affect multiple users and require little or no user interaction to trigger). This underscores the fact that our attack is feasible and poses a privacy risk to numerous users. At the time of writing, we still have ongoing discussions with SteamVR and Unity.
@@
Adoption of Defense. As of August 2024, SteamVR and Rec Room have created patches to defend against our proposed attack, following our suggestions of the aforementioned defense. Specifically, in SteamVR Beta 2.7.1 [17], SteamVR restricts VR applications from accessing hand motion data when the Steam keyboard is opened in Steam Overlay, so that the applications cannot inadvertently synchronize the typing motion data with remote users. Similarly, in Rec Room’s “My Little Template” Edition update [14], Rec Room stops synchronizing a user’s hand motion data with remote users when the user is typing in a sensitive field (e.g., password).
@@ Appendix A Example of a Packet Parsed with Generic Photon Protocol Parser
In Figure 13, Figure 14, Figure 15, and Figure 16, we show how the packet payload is parsed by a generic parser for the Photon protocol (Section 4.2). Photon is used to exchange objects between game clients and the Rec Room server. Note that, at this stage of parsing, we (the attacker) do not (yet) know the meaning (semantics) of specific fields of these objects, which are structured in different ways for different applications. However, the Photon parser can decode the raw packet payload into objects.
@@
Figure 14: Example of a parsed motion data packet from Galaxity.
Figure 15: Example of a parsed motion data packet from Sing Together: VR Karaoke.
Figure 16: Example of a parsed motion data packet from oVRshot.
@@ Appendix D Other Results for the Keylogging Attack
Appendix E Additional Applications
We recovered motion data from the network packets collected in the applications listed in Table 4.
App Name Genre
Galaxity social game
Sing Together: VR Karaoke music game
oVRshot shooter game
Table 4: List of additional applications on which we reproduced the steps to extract motion data.
Appendix F More Analysis for Attack on Partially Reverse Engineered Packets